The reaction

SO2(g)+2H2S(g) ⇌ 3 S(s)+2H2O(g)

is the basis of a suggested method for removal of SO2 , a pollutant that irritates airways causing coughing, from power-plant stack gases.

The values below may be helpful when answering questions about the process.

Substance ΔG∘f
(kJ/mol)
ΔH∘f
(kJ/mol)
H2O(g) − 228.6 − 241.8
H2O(l) − 237.1 − 285.8
SO2(g) − 300.4 − 296.9
SO3(g) − 370.4 − 395.2
H2S(g) − 33.01 − 20.17
S(s) 0 0

Assume that the partial pressure of sulfur dioxide, PSO2, is equal to the partial pressure of dihydrogen sulfide, PH2S, and therefore PSO2=PH2S. If the vapor pressure of water is 22 torr , calculate the equilibrium partial pressure of SO2 (PSO2) in the system at 298 K.

Express the pressure in atmospheres to two significant figures.
Answer #1

Calculate ΔG0rxn as below.

ΔG0rxn = ∑n.ΔG0f(products) - ∑n.ΔG0f(reactants)

where n denotes the number of mol(s) of reactant and product molecules. Therefore,

ΔG0rxn = [(3 mol S)*ΔG0f(S, s) + (2 mol H2O)*ΔG0f(H2O, g)] – [(1 mol SO2)*ΔG0f(SO2, g) + (2 mol H2S)*ΔG0f(H2S)]

= [(3 mol)*(0 kJ/mol) + (2 mol)*(-228.6 kJ/mol)] – [(1 mol)*(-300.4 kJ/mol) + (2 mol)*(-33.01 kJ/mol)]

= (-457.2 kJ) – (-300.4 kJ – 66.02 kJ)

= (-457.2 kJ) – (-366.42 kJ)

= -90.78 kJ

Determine the equilibrium constant for the reaction as

ΔG0rxn = -RTln Kp

where T = 298 K is the temperature of the reaction and Kp is the equilibrium constant for the reaction.

Plug in values and obtain

-90.78 kJ = -(8.314 J/mol.K)*(298 K)*ln Kp

======> (-90.78 kJ)*(1000 J)/(1 kJ) = -(2477.572 J/mol )*ln Kp

======> -9078 J = -(2477.572 J/mol)*ln Kp

======> ln Kp =(-9078 J)/(-2477.572 J/mol)

======> ln Kp = 3.6641 (ignore units)

======> Kp = exp(3.6641)

======> Kp = 39.02

Write down the expression for Kp as

Kp = (PH2O)2/(PSO2)(PH2S)2

Given the relation and PH2O = 22 torr (ignore units), we can write

39.02 = (22)2/(P)(P)2 (denote PSO2 = PH2S = P)

=====> 39.02 = (484)/(P)3

=====> P3 = 484/39.02 = 12.40

=====> P = 2.314 ≈ 2.3 (correct to 2 sig. figs)

The partial pressure of SO2 is 2.3 torr (ans).